# MMSci: A Dataset for Graduate-Level Multi-Discipline Multimodal Scientific Understanding

### v2

@@ 8 Conclusion
In this work, we present MMSci, a multidisciplinary multimodal dataset containing high-quality, peer-reviewed articles and figures across 72 scientific disciplines. Using this dataset, we construct a challenging benchmark to evaluate the capabilities of MLLMs in understanding scientific figures and content, revealing significant deficiencies. Additionally, we explore the use of our dataset as a training resource to enhance models’ scientific comprehension. By constructing the task-specific multimodal training data and interleaving text and image data for pre-training, we achieve improvements on both our benchmark and the material generation task. Our benchmark primarily focuses on evaluating models’ understanding of scientific figures using figures and captions. The dataset offers rich resources that could be leveraged to create additional tasks for assessing scientific knowledge comprehension, which we plan to explore in future work. Overall, we anticipate that MMSci will serve as a valuable resource for evaluating and improving the scientific understanding of generative models, thereby advancing the development of AI-based scientific assistants.
Ethics Statements
We collected the dataset by crawling articles published in Nature Communications, which are freely and permanently available online upon publication, with no subscription fees or registration barriers [^6] https://www.nature.com/ncomms/open-access. These articles are licensed under the Creative Commons Attribution 4.0 International (CC BY) license. The potential risks or misuse of the dataset include: (1) Misuse in Academic Integrity: The use of AI research assistants based on this dataset may lead to issues such as academic fraud, data fabrication, or improper assistance in academic work. We strongly advise users to employ these tools responsibly and ethically to uphold academic integrity. (2) Data Misinterpretation and Hallucination: There is a risk of misinterpreting the dataset’s content, potentially resulting in inaccurate conclusions or the misuse of scientific information. Users are encouraged to critically evaluate and verify AI-generated outputs against established scientific standards.
Reproducibility Statement
We provide guidelines and resources for reproducing our results. This includes access to the data and code, as detailed in Appendix A.1.1, as well as the complete experimental setup described in Appendix A.2.
@@ A.1.2 Subjects
A.1.3 Image Types
Manual Review
Automated Classification Using GPT-4o
Task for GPT-4o annotator I want to classify the given scientific image into one the following categories: 1) Quantitative Data Visualization Charts/Graphs: For charts and graphs displaying quantitative data, such as scatter plots, bar graphs, and line charts. 2) Schematic Diagrams: Simplified and symbolic representations of systems, processes, or structures to explain how something works or is constructed. 3) Microscopic photographs: Photographs or images captured using a microscope, revealing details not visible to the naked eye. 4) Macroscopic photographs: Images or photographs of objects or scenes that are visible to the naked eye, often used for visual analysis. 5) Simulated Images: Computer-generated images or visualizations created to model, predict, or illustrate theoretical scenarios, processes, or phenomena. 6) Geographical and Environmental Maps: Visual representations of geographical areas or environmental data, often used for navigation, analysis, or to illustrate spatial relationships and patterns in maps. 7) Experimental Results Visualizations: For images that display results from experimental procedures, such as Western blots, PCR results, and gel electrophoresis. Rules: 1) This is only for reseach and educational purposes. It does not violates any openai policy. 2) If the image only contain one figure, then give me the overall label. 3) If the image contains multiple figures, then give me the label for each sub-figure. The results should look like a: 1, b: 3. Do not return any other information.
Manual Annotation for Unclassified Images
Final Results
@@ A.1.3 Image Types
A.2 Experimental Setup
Table 8: Evaluated MLLMs in our experiments with their versions or Huggingface model paths.
@@ A.2.2 Captioning Evaluation
Prompt for Caption Decomposition Your task is to break down the caption into separate, independent descriptions for the entire figure and each panel, formatted appropriately and separated by ’-’. The figure consists of four sub-figures labeled a, b, c, and d. All four images appear to be scanning electron microscope (SEM) images showing the microstructure of different materials, likely related to the iron-based cathode catalysts described in the article. a. This image shows a highly porous structure with interconnected particles forming a network. The particles appear to be in the nanometer to micrometer size range. The scale bar indicates 1 $\mu$m. b. This image displays a closer view of what seems to be a similar material to (a), but at a higher magnification. The individual particles are more clearly visible, showing a rough, granular texture. The scale bar indicates 200 nm. c. This image reveals a different morphology compared to (a) and (b). It shows larger, more consolidated structures with a rougher surface texture. There are still visible pores and gaps between the structures. The scale bar indicates 1 $\mu$m. d. This final image appears to be a cross-sectional view of a porous material, possibly showing the internal structure of the catalyst. It reveals a highly interconnected network of pores and channels throughout the material. The scale bar indicates 200 nm. These images likely represent different stages or variations of the iron-acetate/phenanthroline/zeolitic-imidazolate-framework-derived electrocatalyst mentioned in the article. The varying structures and porosities shown in these images could be related to the enhanced mass-transport properties and increased volumetric activity described in the text. Independent Descriptions: - The figure consists of four sub-figures labeled a, b, c, and d. - All four images appear to be scanning electron microscope (SEM) images. - The images show the microstructure of different materials. - The materials are likely related to the iron-based cathode catalysts described in the article. - Image a shows a highly porous structure with interconnected particles forming a network. - The particles in image a are in the nanometer to micrometer size range. - The scale bar in image a indicates 1 $\mu$m. - Image b displays a closer view of a material similar to that in image a but at higher magnification. - The individual particles in image b are more clearly visible and show a rough, granular texture. - The scale bar in image b indicates 200 nm. - Image c shows larger, more consolidated structures with a rougher surface texture. - There are visible pores and gaps between the structures in image c. - The scale bar in image c indicates 1 $\mu$m. - Image d appears to be a cross-sectional view of a porous material. - Image d reveals the internal structure of the catalyst with a highly interconnected network of pores and channels. - The scale bar in image d indicates 200 nm. - These images likely represent different stages or variations of the iron-acetate/phenanthroline/zeolitic-imidazolate-framework-derived electrocatalyst mentioned in the article. - The varying structures and porosities shown in these images could be related to the enhanced mass-transport properties described in the text. - The varying structures and porosities in the images may contribute to increased volumetric activity described in the article.
@@
Prompt for Atom Unit Description Rating How relevant is the generated caption to the provided human-written caption for the figure? Determine the extent to which the information in the generated caption is included or referenced in the human-written caption. Respond with a score between 0 and 5. Human-written caption: {REFERENCE} Generated caption: {GENERATION}
@@
Prompt for G-Eval Evaluation You will be given a oracle caption that describes a figure. You will then be given a second caption written for the same figure. Your task is to rate the second caption on one the following metric. Evaluation Criteria: Relevance (1-5) - The extent to which the second caption is relevant to the key elements and context described in the oracle caption. A relevant caption should focus on the same subjects, objects, actions, or context highlighted in the oracle caption, without introducing unrelated or extraneous details. Evaluation Steps: 1. Review the Oracle Caption: Carefully read the oracle caption to understand the main elements and context it describes. 2. Review the Second Caption: Assess whether the second caption focuses on the same key elements and context as the oracle caption. Evaluate if the second caption stays on topic and does not introduce irrelevant details. 3. Assign a Score for Relevance: Based on the Evaluation Criteria, rate how relevant the second caption is to the oracle caption’s description of the same image.
@@
base model https://huggingface.co/Qwen/Qwen2-VL-2B-Instruct
@@
A.2.3 Visual Supervised Fine-tuning
@@
A.2.4 Visual Language Pre-training
@@ A.3.2 Intended Use
This dataset is used to evaluate and enhance the large multimodal models (MLLMs)’ understanding of advanced multimodal scientific knowledge.
@@ A.3.4 Social Impact and Ethical Considerations
The benefits of our dataset are two-fold: (1) Evaluation Benchmark: This dataset serves as a valuable evaluation benchmark for assessing the understanding of large multimodal models (MLLMs) regarding scientific articles and figures. (2) Training Resources: It can be used as a training resource to enhance MLLMs’ understanding of scientific articles and figures, improving their performance in various scientific and research-related tasks.
@@ A.3.5 Limitations
Our dataset MMSci provides a comprehensive multimodal dataset across 72 scientific disciplines and serves as both a benchmark and a training resource. However, there are some limitations in our current exploration. (1) Due to limited resources, we were unable to evaluate a wide range of large-scale open-source MLLMs. (2) Our benchmark primarily assesses models’ understanding of scientific figures using the figures and captions. The dataset still provide other valuable resources that could be used to create additional tasks, such as single- and multimodal questions aimed at evaluating models’ scientific knowledge. We plan to explore these opportunities in future work. Despite these limitations, we believe MMSci will be a valuable resource for the research community. All data will be made publicly available.
@@ A.4 Examples
We present several figures as our case study to illustrate multiple-choice questions under three setting in Figure 8, 9, 10, respectively.

### v3

@@ 8 Conclusion
In this work, we present MMSci, a multidisciplinary multimodal dataset containing high-quality, peer-reviewed articles and figures across 72 scientific disciplines. Using this dataset, we construct a challenging benchmark to evaluate the capabilities of LVLMs in understanding scientific figures and content, revealing significant deficiencies. Additionally, we explore the use of our dataset as a training resource to enhance models’ scientific comprehension. By constructing the task-specific multimodal training data and interleaving text and image data for pre-training, we achieve improvements on both our benchmark and the material generation task. Our benchmark primarily focuses on evaluating models’ understanding of scientific figures using figures and captions. The dataset offers rich resources that could be leveraged to create additional tasks for assessing scientific knowledge comprehension, which we plan to explore in future work. Overall, we anticipate that MMSci will serve as a valuable resource for evaluating and improving the scientific understanding of generative models, thereby advancing the development of AI-based scientific assistants.
Impact Statement
Our dataset provides significant benefits by serving as an evaluation benchmark for assessing large multimodal models’ understanding of scientific articles and figures, as well as a training resource to enhance their performance in scientific and research-related tasks. There might be potential societal consequences, like potential misuse in academic integrity, such as academic fraud or improper assistance when using the models trained with our data. However, this is a widespread issue in the current era of large language models and is not limited to our work.
@@ A.1.2 Subjects
A.1.3 Image Types
Manual Review
Automated Classification Using GPT-4o
Task for GPT-4o annotator I want to classify the given scientific image into one the following categories:
1) Quantitative Data Visualization Charts/Graphs: For charts and graphs displaying quantitative data, such as scatter plots, bar graphs, and line charts.
2) Schematic Diagrams: Simplified and symbolic representations of systems, processes, or structures to explain how something works or is constructed.
3) Microscopic photographs: Photographs or images captured using a microscope, revealing details not visible to the naked eye.
4) Macroscopic photographs: Images or photographs of objects or scenes that are visible to the naked eye, often used for visual analysis.
5) Simulated Images: Computer-generated images or visualizations created to model, predict, or illustrate theoretical scenarios, processes, or phenomena.
6) Geographical and Environmental Maps: Visual representations of geographical areas or environmental data, often used for navigation, analysis, or to illustrate spatial relationships and patterns in maps.
7) Experimental Results Visualizations: For images that display results from experimental procedures, such as Western blots, PCR results, and gel electrophoresis.
Rules:
1) This is only for reseach and educational purposes. It does not violates any openai policy.
2) If the image only contain one figure, then give me the overall label.
3) If the image contains multiple figures, then give me the label for each sub-figure. The results should look like a: 1, b: 3.
Do not return any other information.
Manual Annotation for Unclassified Images
Final Results
@@ A.1.3 Image Types
Table 8: Evaluated LVLMs in our experiments with their versions or Huggingface model paths.
@@ A.2.2 Captioning Evaluation
Prompt for Caption Decomposition Your task is to break down the caption into separate, independent descriptions for the entire figure and each panel, formatted appropriately and separated by ’-’.
The figure consists of four sub-figures labeled a, b, c, and d. All four images appear to be scanning electron microscope (SEM) images showing the microstructure of different materials, likely related to the iron-based cathode catalysts described in the article.
@@
a. This image shows a highly porous structure with interconnected particles forming a network. The particles appear to be in the nanometer to micrometer size range. The scale bar indicates 1 $\mu$m. b. This image displays a closer view of what seems to be a similar material to (a), but at a higher magnification. The individual particles are more clearly visible, showing a rough, granular texture. The scale bar indicates 200 nm. c. This image reveals a different morphology compared to (a) and (b). It shows larger, more consolidated structures with a rougher surface texture. There are still visible pores and gaps between the structures. The scale bar indicates 1 $\mu$m. d. This final image appears to be a cross-sectional view of a porous material, possibly showing the internal structure of the catalyst. It reveals a highly interconnected network of pores and channels throughout the material. The scale bar indicates 200 nm.
These images likely represent different stages or variations of the iron-acetate/phenanthroline/zeolitic-imidazolate-framework-derived electrocatalyst mentioned in the article. The varying structures and porosities shown in these images could be related to the enhanced mass-transport properties and increased volumetric activity described in the text.
Independent Descriptions:
- The figure consists of four sub-figures labeled a, b, c, and d.
- All four images appear to be scanning electron microscope (SEM) images.
- The images show the microstructure of different materials.
- The materials are likely related to the iron-based cathode catalysts described in the article.
- Image a shows a highly porous structure with interconnected particles forming a network.
- The particles in image a are in the nanometer to micrometer size range. - The scale bar in image a indicates 1 $\mu$m.
- Image b displays a closer view of a material similar to that in image a but at higher magnification.
- The individual particles in image b are more clearly visible and show a rough, granular texture.
- The scale bar in image b indicates 200 nm.
- Image c shows larger, more consolidated structures with a rougher surface texture.
- There are visible pores and gaps between the structures in image c.
- The scale bar in image c indicates 1 $\mu$m.
- Image d appears to be a cross-sectional view of a porous material.
- Image d reveals the internal structure of the catalyst with a highly interconnected network of pores and channels.
- The scale bar in image d indicates 200 nm.
- These images likely represent different stages or variations of the iron-acetate/phenanthroline/zeolitic-imidazolate-framework-derived electrocatalyst mentioned in the article.
- The varying structures and porosities shown in these images could be related to the enhanced mass-transport properties described in the text.
- The varying structures and porosities in the images may contribute to increased volumetric activity described in the article.
@@
Prompt for Atom Unit Description Rating How relevant is the generated caption to the provided human-written caption for the figure? Determine the extent to which the information in the generated caption is included or referenced in the human-written caption. Respond with a score between 0 and 5.
Human-written caption: {REFERENCE}
Generated caption: {GENERATION}
@@
Prompt for G-Eval Evaluation You will be given a oracle caption that describes a figure. You will then be given a second caption written for the same figure. Your task is to rate the second caption on one the following metric.
Evaluation Criteria: Relevance (1-5) - The extent to which the second caption is relevant to the key elements and context described in the oracle caption. A relevant caption should focus on the same subjects, objects, actions, or context highlighted in the oracle caption, without introducing unrelated or extraneous details.
Evaluation Steps: 1. Review the Oracle Caption: Carefully read the oracle caption to understand the main elements and context it describes.
2. Review the Second Caption: Assess whether the second caption focuses on the same key elements and context as the oracle caption. Evaluate if the second caption stays on topic and does not introduce irrelevant details.
3. Assign a Score for Relevance: Based on the Evaluation Criteria, rate how relevant the second caption is to the oracle caption’s description of the same image.
@@
base model https://huggingface.co/Qwen/Qwen2-VL-7B-Instruct
@@
Table 10: Performance comparison on scientific figure captioning task when grounded on abstract and full article.
Model Context BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L Meteor BertScore FActScore G-Eval
Gemini-1.5-Flash Abstract 3.29 26.74 7.47 16.03 28.71 81.80 10.14 4.08
Gemini-1.5-Pro Abstract 3.33 28.71 7.73 16.89 28.91 81.93 13.76 4.08
Claude-3.5-Sonnet Abstract 3.20 29.60 6.71 16.65 27.52 81.76 12.11 4.04
GPT-4V Abstract 3.18 28.45 7.01 15.65 27.62 82.37 19.52 4.13
GPT-4o Abstract 3.58 28.85 7.79 16.36 28.37 81.84 18.87 4.22
Gemini-1.5-Flash Full Article 6.94 32.83 14.15 22.02 34.50 83.26 19.41 4.12
Gemini-1.5-Pro Full Article 7.01 32.24 13.34 19.32 33.75 83.18 19.33 4.22
Claude-3.5-Sonnet Full Article 7.99 37.63 13.61 23.63 34.66 84.34 21.67 4.52
GPT-4V Full Article 5.65 33.09 10.95 19.25 31.46 83.48 23.18 4.24
GPT-4o Full Article 9.90 37.06 17.63 24.89 37.52 83.64 24.12 4.58
Captioning Grounded on Full Article
We also explored using entire articles as context for captioning. Due to the average article length exceeding 10k tokens, we evaluated this approach only on proprietary models capable of handling long contexts: GPT-4o, GPT-4V, Claude-3.5-Sonnet, and Gemini-1.5-Pro/Flash. As shown in Table 10, providing the full article as context improved performance compared to using only the abstract. This improvement is reasonable since understanding scientific figures typically requires grounding in the article’s content, as abstracts alone may not provide sufficient context. However, we note that this approach may potentially benefit from content repetition, as similar descriptions might appear in both the caption and the article text.
Table 11: Recategorization of the 72 subjects in MMSci dataset for recruiting Phd experts of each major category from Prolific platform.
Re-categoried Fields Original Subjects from Nature Communications
Material Science Materials science, Nanoscience and technology
Chemistry Chemistry
Physics Physics, Optics and photonics
Engineering Engineering
Energy Energy science and technology, Energy and society
Mathematics and Computing Mathematics and computing
Astronomy and Planetary Science Astronomy and planetary science, Planetary science, Space physics
Environment Ecology, Environmental sciences, Biogeochemistry, Water resources
Climate Sciences Climate sciences
Earth Solid Earth sciences, Ocean sciences, Natural hazards, Hydrology, Limnology, Geography
Social Sciences Environmental social sciences, Psychology, Social sciences, Scientific community, Developing world
Biochemistry Biochemistry, Molecular biology, Biophysics, Structural biology, Chemical biology
Biological Sciences Microbiology, Genetics, Biological techniques, Computational biology and bioinformatics, Developmental biology, Evolution, Plant sciences, Physiology, Systems biology, Zoology, Cell biology
Biomedical Sciences Neuroscience, Immunology, Biotechnology, Stem cells, Pathogenesis, Biomarkers, Anatomy, Molecular medicine
Health and Medicine Cancer, Diseases, Medical research, Health care, Oncology, Cardiology, Gastroenterology, Endocrinology, Neurology, Risk factors, Rheumatology, Nephrology, Signs and symptoms, Urology, Health occupations
Pharmacology Drug discovery
Agriculture Agriculture, Forestry
Business and Industry Business and industry
A.2.3 Human Expert Evaluation
To analyze our dataset, we recruited domain experts (PhDs in corresponding fields) through the online professional annotation platform Prolific [^6] https://www.prolific.com/. We refined and consolidated the original 72 subject categories from Prolific into 18 broader groups to balance between comprehensive coverage and sufficient specificity. The recategorized subjects are shown in Table 11. From these 18 recategorized fields, we focused on 10 major scientific domains where PhD annotators were available on Prolific. We recruited 30 PhDs as human evaluators with verified degrees in these domains: Material Science, Chemistry, Physics, Biochemistry, Environment, Climate Sciences, Earth Sciences, Biological Sciences, Biomedical Sciences, and Health and Medicine. Each evaluator provided two types of assessments: Question Quality Assessment and Expert Performance Score. The results for each group are detailed in Table 12.
Question Quality Assessment
For the quality assessment, evaluators were asked to assess whether the questions were clear and demonstrated understanding of scientific knowledge within their respective disciplines. They used the following 5-point scale:
  • Score Point 1: The question is irrelevant or cannot be answered based on the scientific content presented in the figure.
  • Score Point 2: The question lacks clarity or can be answered without specific knowledge of the scientific content in the figure (e.g., it can be answered with common sense).
  • Score Point 3: The question is clear but requires only minimal understanding of the scientific content in the figure.
  • Score Point 4: The question is clear, answerable, and requires an adequate understanding of the scientific content in the figure.
  • Score Point 5: The question is clear, answerable, and effectively evaluates a very deep understanding of the scientific content in the figure.
Expert Performance Score
For the expert evaluation tasks, we created a subset of questions for each category by selecting 25 questions per setting (75 total) from our three figure-caption matching tasks in the original test set. We report the results from the best-performing expert (who achieved the highest average performance) in each category. The annotators were instructed to select their answers within a one-minute time limit per question.
Table 12: Quality scores and Phd experts’ accuracies across the ten re-grouped fields.
Field Fig2Cap SubFig2Cap SubCap2Fig
Quality (1-5) Accuracies (%) Quality (1-5) Accuracies (%) Quality (1-5) Accuracies (%)
Material Science 4.0267 92.00 4.2933 92.00 4.1333 84.00
Chemistry 4.1333 84.00 3.7467 92.00 3.6133 100.00
Physics 4.0267 48.00 3.5467 72.00 3.8133 80.00
Biochemistry 3.1600 80.00 4.8267 56.00 4.4133 72.00
Environment 4.1067 44.00 4.4667 64.00 4.3467 76.00
Climate Sciences 4.1296 77.78 3.6471 88.24 3.4118 82.35
Earth 4.0267 44.00 4.2319 52.17 4.1739 60.87
Biological Sciences 3.8800 48.00 3.6800 48.00 3.7867 32.00
Biomedical Sciences 4.0133 68.00 4.1333 72.00 3.7733 88.00
Health and Medicine 4.3733 56.00 3.7467 80.00 3.6800 52.00
Average 4.0873 64.18 4.0319 71.64 3.9149 72.72
A.2.4 Visual Supervised Fine-tuning
@@
A.2.5 Visual Language Pre-training
@@ A.3.2 Intended Use
This dataset is used to evaluate and enhance the large multimodal models (LVLMs)’ understanding of advanced multimodal scientific knowledge.
@@ A.3.4 Social Impact and Ethical Considerations
The benefits of our dataset are two-fold: (1) Evaluation Benchmark: This dataset serves as a valuable evaluation benchmark for assessing the understanding of large multimodal models (LVLMs) regarding scientific articles and figures. (2) Training Resources: It can be used as a training resource to enhance LVLMs’ understanding of scientific articles and figures, improving their performance in various scientific and research-related tasks.
@@ A.3.5 Limitations
Our dataset MMSci provides a comprehensive multimodal dataset across 72 scientific disciplines and serves as both a benchmark and a training resource. However, there are some limitations in our current exploration. (1) Due to limited resources, we were unable to evaluate a wide range of large-scale open-source LVLMs. (2) Our benchmark primarily assesses models’ understanding of scientific figures using the figures and captions. The dataset still provide other valuable resources that could be used to create additional tasks, such as single- and multimodal questions aimed at evaluating models’ scientific knowledge. We plan to explore these opportunities in future work. Despite these limitations, we believe MMSci will be a valuable resource for the research community. All data will be made publicly available.
@@ A.4 Examples
We present several figures as our case study to illustrate multiple-choice questions under three setting in Figure 7, 8, 9, respectively.